Ms. Green is planting a quarter circle garden in her sqare yard. One side of the yard is 30 feet long . What part of her yard will not be included in the garden ? Give your answer to the nearest tenth of a sqare yard .
Answer by stanbon(75887)

Ms. Green is planting a quarter circle garden in her sqare yard. One side of the yard is 30 feet long . What part of her yard will not be included in the garden ? Give your answer to the nearest tenth of a sqare yard .
------------------
Draw the picture.
You have a circle inscribed in a 10x10 yds.
----
diameter of the circle = 10 yrd
---
Area of the entire circle = pi*10^2 = 100pi sq yds
----
Area not used for garden = (3/4)(100pi) = 75 sq yds
===============
Cheers,
Stan H.
---------------
------